# A Study of Fairness Concerns in AI-based Mobile App Reviews

### v2

@@ 4.2.1. Performance of the classifiers
Table 1 shows the performance of the binary classifiers in detecting fairness reviews. Given that we used 10-fold cross-validation, the score in each row, except the last row (i.e., GPT-3), is the average over the ten folds. Among the classifiers, RoBERTa achieves the highest F1-score (0.82), followed by BERT with an F1-score of 0.80. Considering the accuracy metric, SVM with the combination of TF-IDF and USE as the feature extraction technique reaches the highest accuracy of 0.83, meaning that 83% of predictions (either fairness or non-fairness) are correct. As shown in Table 1, several classifiers such as RoBERTa, Logistic Regression, BERT, and XGBoost reach an accuracy of above 0.80. We see while GPT-3 achieves the highest recall of 0.81, it has low precision of 0.63, meaning that it produces low false negatives and relatively high false positives (more non-fairness reviews are labelled as fairness). RoBERTa with a recall of 0.78 and BERT with a recall of 0.77 are the second-best and third-best classifiers, respectively, from the recall perspective.
@@
Table 1 indicates that Logistic Regression with TF-IDF has higher precision (0.88) than all other classifiers but quite low recall (0.49). Using the combination of three feature extraction techniques (TF-IDF, USE, and Word2Vec) leads to increasing the recall of Logistic Regression from 0.49 to 0.70 and having a precision of 0.85. RoBERTa has also a high precision of 0.85 and a reasonable recall of 0.78. Considering the AUC scores, Logistic Regression with TF-IDF, USE, and Word2Vec performs better than other classifiers.
@@
Given that our ultimate goal is to identify different types of fairness concerns from the user perspective, we decided to select Logistic Regression with TF-IDF, USE, and Word2Vec as our best binary classifier (see Table 1). It is because (1) on average, this model has the highest AUC, high precision and good F1-score. (2) Logistic Regression with TF-IDF, USE, and Word2Vec reaches a precision of 0.94, AUC of 0.93, and recall of 0.77 in the seventh fold. In other words, inspired by the methodology used in (Nema et al., 2022), we prioritized AUC and precision over recall. We used this model that reaches a precision of 0.94 in the seventh fold for the rest of our work (Section 4.2.2, Section 4.2.3, and Section 5). Particularly, this model could minimize the number of noises in our clusters (see Section 5).
@@ 4.2.2. Distribution of fairness reviews across app categories
We were also interested in knowing the distribution of fairness concerns in different app categories. Remember that in Section 3.1.2, we collected 9,475,506 reviews from 23 distinct app categories. As shown in Fig. 3, we applied our best classifier (Logistic Regression with TF-IDF, USE, and Word2Vec) on this number of reviews and found 91,974 fairness reviews (0.97% of all reviews include at least one discussion around fairness). Table 2 shows the number of reviews and the percentage of fairness reviews in 23 app categories. As shown in Table 2, the ‘communication’ category has the highest percentage of fairness reviews (1.76%), followed by the ‘social’ category (1.64%) and ‘entertainment’ category (1.45%). We also found that app users in the ‘personalization’ and ‘auto & vehicles’ categories rarely discuss fairness issues in reviews (i.e., less than 0.1% of their reviews are about fairness).
Figure 3. Applying the best-binary classifier and K-means
@@ 4.2.3. Patterns in the extracted fairness reviews
Our analysis shows that the binary classifiers, particularly the best-performing model, can identify reviews that do not include any of the 21 fairness-related keywords collected in Section 3.1. For instance, this model correctly classified the following reviews as fairness reviews: “I love the app but I need a dark mode. I’m on Android and I know that iOS has a dark mode option available. Dark mode would make this 5 stars.” and “[The app is] not good in the language. [It] does not support the Arabic language”. In the first review, the user seemingly criticizes the app provider as it does not provide the same dark mode feature that is available for iOS users to Android users. The second review complains about the app as it does not support the Arabic language. In both examples, such unfair behaviours may lead to the exclusion of or discrimination towards some users. We also found some patterns in the reviews misclassified by our model that can be investigated in future to enhance the model. Some of these reviews include one or more of the 21 fairness-related keywords but they do not raise any concrete or specific fairness concerns or discuss root causes of raised fairness concerns. For example, the following review “Easy to use and alerts are prompt! [I] just wish I could get my hands on the tech grey. For people [who] have been loyal customers literally their entire lives, I think it’s unfair not to be able to purchase a product they really want but can’t due to it being sold out.” criticizes a product produced by a company and does not point out any concrete fairness concern regarding the app. We labelled such reviews as non-fairness reviews in our ground-truth dataset. However, our model sometimes classifies such reviews as a fairness review (a false positive review). It is possibly because the model makes the choice based on a fairness-related keyword (e.g., “unfair” in the above review) and the negative sentiment of the review. This may suggest that our model needs a more diverse and larger dataset for training and testing.
@@
All
reviews
Fairness
reviews
% of fairness
reviews
@@
Key Finding 1: Our experiments show that Logistic Regression, combining TF-IDF, USE, and Word2Vec as the feature extraction technique, is the best-performing binary classifier to detect fairness reviews and reaches the highest precision of 94%.
5. Fairness Clustering and Summarization
As discussed in Section 4, the binary classifiers can accurately identify fairness reviews from a large number of reviews. However, they do not provide a fine-grained view of different types of fairness concerns raised by various users of AI-based apps. Hence, we aimed to cluster and summarize fairness reviews, as discussed next.
@@ 5.1.1. Dataset
As discussed in Section 4.2, Logistic Regression was selected as the best-performing classifier because it had the highest precision and AUC. Once we applied it on 9,475,506 reviews, we got 91,974 fairness reviews (see Fig. 3). We used these 91,974 reviews as a dataset for the clustering and summarization approach.
@@ 5.1.2. K-means Clustering
In addition to the summarization metric, we used the silhouette score to determine the quality of clusters (Caliński and Harabasz, 1974). The silhouette score is a measure of how similar a data point is to the other data points in its cluster compared to how similar it is to the data points in other clusters. The silhouette score thus indicates whether a given cluster is well-formed or not. We executed K-means with different $k$ ranging from 2 to 10. For each $k$, K-means was run 100 times with different centroid seeds. We also set the maximum number of iterations for a single run to 100. The final result of K-means is the best output out of 100 runs in terms of inertia, which is the sum of squared distances of samples to their closest cluster center. Our experiments showed that the highest value for the summarization metric is achieved when k=10 (Fig. 3). Fig. 4 presents the distribution of the silhouette scores of reviews in each of the 10 clusters (k=10). The dotted red line shows the average silhouette score for k=10. As shown in Fig. 4, at least 40% of the reviews in clusters 2 to 9 have a silhouette score more than the average (the dotted red line). Clusters 2 to 9 are considered compact clusters. We ignored clusters 0 and 1 because more than 60% of their reviews have a silhouette score below the average.
@@
To understand the main fairness concern discussed in each of the eight compact clusters, we descendingly sorted the reviews in each compact cluster based on their silhouette scores. The first author manually analyzed the 30 reviews with the highest silhouette scores in each cluster ($8*30=240$ reviews in total) and suggested a topic for each cluster. Then, three other authors rechecked all these 240 reviews and the proposed topics. The first author and the rest of authors held several meetings to discuss the topics, refined them and solved any disagreements. Although we had eight compact clusters, as shown in Fig. 3, the qualitative analysis process led to the identification of six distinct fairness concerns that are discussed next in Section 5.2.
@@ 5.2. Results
6. Root Causes of Fairness Concerns
@@ 6.1. Approach
As shown in Fig. 3, our binary classifier identified 91,974 fairness reviews (Section 4.2), which were clustered into six types of fairness concerns (Section 5.2). These 91,974 app reviews received 4,911 responses from app owners. To understand the root causes reported by app owners for the identified fairness concerns, we randomly selected 2,248 out of the 4,911 app owners’ responses to fairness reviews with a confidence level of 99% and margin error of 2% (Kadam and Bhalerao, 2010). In the first step, the first author conducted the open coding procedure (Glaser and Strauss, 2017) on these 2,248 app owners’ responses. Next, the third author reviewed all the codes and categories produced by the first author. Last, the first and third authors held several meetings to discuss and refine the codes and categories. Similar to Section 3.1, the negotiated agreement approach (Campbell et al., 2013) was employed to resolve any disagreements and conflicts between the authors. This process led to the identification of six root causes of fairness concerns.
@@ 6.2. Results
1. Copyright Issues. App owners admit that some of the fairness concerns are due to the copyright issues imposed on the apps, platforms, or devices. This can mainly be a reason behind Fairness Concern 1 (receiving different quality of features and services in different platforms and devices) identified in RQ1 (Section 5.2). For example in the following conversation, the copyright issue is mentioned by an app owner as the reason why Android and iOS users do not receive the same emojis.
@@
3. Buggy Code. Some app owners admit that buggy code can be a reason behind the identified fairness concerns (particularly Fairness Concerns 4 and 5).
@@ 7. Discussion
Benefits of findings for mobile app development tasks. In a systematic review, Dąbrowski et al. (Dąbrowski et al., 2022) found that app review analysis can support various software development activities (e.g., requirement engineering and design). We argue that the findings obtained in RQ1 (fairness concerns from the users’ perceptive) and RQ2 (root causes of fairness concerns from the app owners’ viewpoint) could benefit mobile app development in different ways. (1) The identified fairness concerns and their potential root causes can help app owners and developers understand and elicit users’ fairness-related requirements and expectations. For example, the identified fairness concerns can guide app developers on questions that should be asked of users to gather fairness-related requirements and expectations. (2) The knowledge of potential fairness concerns can help app developers be aware of and avoid common fairness concerns during app development. (3) The identified fairness concerns can act as a checklist for app developers to enable them to specify test scenarios that evaluate specific fairness concerns. (4) Finally, our findings can inform researchers and tool builders to develop tools to embed fairness in AI-based apps, and detect and fix fairness concerns.
@@ 8. Threats to Validity
Internal Validity. A key concern in our internal validity is the inherent ambiguity attached to the concept of ‘fairness’, which can be nuanced, with different interpretations. App users might express fairness concerns using various terms and phrases, not all of which are easy to capture by a keyword-based approach like ours. To mitigate the potential impact of this threat, we used an iterative and multi-faceted approach to identify fairness-related keywords. We started with a literature-driven set of keywords and expanded this using KeyBERT. The keyword set was reviewed and refined manually to ensure relevance and accuracy. Moreover, our manual labelling process in RQ1 and qualitative analysis in RQ2 included multiple coders to minimize individual bias and capture a broader interpretation of fairness.
Conclusion Validity. The conclusions regarding the fairness concerns of users and the root causes of the fairness concerns were based on the results of our clustering approach. A different instrumentation of our clustering approach might have influenced our conclusions in Section 5.2 and Section 6.2. To mitigate the potential impact of this threat, we strived for transparency and thoroughness in our reporting. We explained our choices in the clustering process, including why a particular technique was selected and our parameter-tuning process. We also acknowledge the influence of these choices on our findings. While our choices were reasonable and based on best practices, further experimentation with more techniques is required in our future work.
External Validity. Our evaluation is based on 108 apps from 23 categories and an overall analysis of $\approx$9.5 million reviews. The results we received across these apps with a thorough manual analysis of $>$2,600 reviews for building our classifier, analysis of $>$92,000 reviews using the classifier and clustering, and qualitative analysis of 2,248 lend a degree of confidence to the generalizability of our study. Further experimentation is nonetheless required to mitigate external validity threats.

### v3

@@ 4.2.1. Performance of the models
Table 3 shows the performance of the models (described in Section 4.1.2) in detecting fairness reviews. Given that we used 10-fold cross-validation for ML and pre-trained models, the score in each row, except the last row (i.e., GPT-3), is the average over the ten folds. Among the models, RoBERTa achieves the highest F1-score (0.82), followed by BERT with an F1-score of 0.80. Considering the accuracy metric, SVM with the combination of TF-IDF and USE as the feature extraction technique reaches the highest accuracy of 0.83, meaning that 83% of predictions (either fairness or non-fairness) are correct. As shown in Table 3, several models such as RoBERTa, Logistic Regression, BERT, and XGBoost reach an accuracy of above 0.80. We see while GPT-3 achieves the highest recall of 0.81, it has low precision of 0.63, meaning that it produces low false negatives and relatively high false positives (more non-fairness reviews are labeled as fairness). RoBERTa with a recall of 0.78 and BERT with a recall of 0.77 are the second-best and third-best models, respectively, from the recall perspective.
@@
Table 3 indicates that Logistic Regression with TF-IDF has higher precision (0.88) than all other models but quite low recall (0.49). Using the combination of three feature extraction techniques (TF-IDF, USE, and Word2Vec) leads to increasing the recall of Logistic Regression from 0.49 to 0.70 and having a precision of 0.85. RoBERTa has also a high precision of 0.85 and a reasonable recall of 0.78. Considering the AUC scores, Logistic Regression with TF-IDF, USE, and Word2Vec performs better than other models.
@@
Given that our ultimate goal is to identify different types of fairness concerns from the user perspective, we decided to select Logistic Regression with TF-IDF, USE, and Word2Vec as our best model (see Table 3). It is because (1) on average, this model has the highest AUC, high precision and good F1-score. (2) Logistic Regression with TF-IDF, USE, and Word2Vec reaches a precision of 0.94, AUC of 0.93, and recall of 0.77 in the seventh fold. In other words, inspired by the methodology used in (Nema et al., 2022), we prioritized AUC and precision over recall. We used this model that reaches a precision of 0.94 in the seventh fold for the rest of our work (Section 4.2.3, Section 4.2.4, and Section 5). Particularly, this model could minimize the number of noises in our clusters (see Section 5).
Table 4. Results of best-performing model on a sample of all reviews dataset
P R F1 ACC AUC
LR+TDIDF+USE+W2V 88.75 95.17 91.85 92.12 92.32
Figure 3. Confusion matrix of our best performing model for all reviews
4.2.2. Performance of the best-performing model on all reviews
As discussed in Section 3.1.2, we collected 9,475,506 reviews from 23 distinct app categories. As shown in Fig. 4, we applied our best model (Logistic Regression with TF-IDF, USE, and Word2Vec) on this number of reviews. The best model classified the 9,475,506 reviews into 91,974 fairness reviews and 9,383,532 non-fairness reviews. To evaluate the performance of the best model on all reviews, we randomly selected 400 reviews from 91,974 non-fairness reviews and 400 reviews from 91,974 fairness reviews with a confidence level of 95% and a margin of error of 5% (Israel et al., 1992; Ahmad and Halim, 2017). The first and third authors undertook the task of independently labeling these 800 reviews. The inter-rater agreement between them was a Cohen’s kappa coefficient of 0.92, indicating perfect agreement. They had 30 disagreements, which were resolved through the negotiated agreement approach (Campbell et al., 2013). Table 4 shows that applying the best model on all reviews yielded good results in all metrics: a precision of 88.75%, a recall of 95.17%, an F1 of 91.85%, an ACC of 92.12%, an AUC of 92.32%. Fig. 3 also shows the confusion matrix of the best model for all reviews.
@@ 4.2.3. Distribution of fairness reviews across app categories
We also wanted to know the distribution of fairness reviews in different app categories. Section 4.2.2 shows that 91,974 reviews (0.97%) out of all reviews (9,475,506) were classified as fairness reviews. Table 5 indicates the number of reviews and the percentage of fairness reviews in each app category. As shown in Table 5, the ‘communication’ category has the highest percentage of fairness reviews (1.76%), followed by the ‘social’ category (1.64%) and ‘entertainment’ category (1.45%). We also found that app users in the ‘personalization’ and ‘auto & vehicles’ categories rarely discuss fairness issues in reviews (i.e., less than 0.1% of their reviews are about fairness).
Figure 4. Applying the best model and K-means
@@ 4.2.4. Patterns in the extracted fairness reviews
Our analysis shows that the models, particularly the best-performing model, can identify reviews that do not include any of the 21 fairness-related keywords collected in Section 3.1. For instance, this model correctly classified the following reviews as fairness reviews: “I love the app but I need a dark mode. I’m on Android and I know that iOS has a dark mode option available. Dark mode would make this 5 stars.” and “[The app is] not good in the language. [It] does not support the Arabic language”. In the first review, the user seemingly criticizes the app provider as it does not provide the same dark mode feature that is available for iOS users to Android users. The second review complains about the app as it does not support the Arabic language. In both examples, such unfair behaviours may lead to the exclusion of or discrimination towards some users. We also found some patterns in the reviews misclassified by our model that can be investigated in future to enhance the model. Some of these reviews include one or more of the 21 fairness-related keywords but they do not raise any concrete or specific fairness concerns or discuss root causes of raised fairness concerns. For example, the following review “Easy to use and alerts are prompt! [I] just wish I could get my hands on the tech grey. For people [who] have been loyal customers literally their entire lives, I think it’s unfair not to be able to purchase a product they really want but can’t due to it being sold out.” criticizes a product produced by a company and does not point out any concrete fairness concern regarding the app. We labeled such reviews as non-fairness reviews in our ground-truth dataset. However, our model sometimes classifies such reviews as a fairness review (a false positive review). It is possibly because the model makes the choice based on a fairness-related keyword (e.g., “unfair” in the above review) and the negative sentiment of the review. This may suggest that our model needs a more diverse and larger dataset for training and testing.
@@
All reviews
Fairness reviews
% of fairness reviews
@@
Key Finding 1: Our experiments show that Logistic Regression, combining TF-IDF, USE, and Word2Vec as the feature extraction technique, is the best-performing model to detect fairness reviews and reaches the highest precision of 94%.
5. Fairness Clustering and Summarization (RQ2)
As discussed in Section 4, the models can accurately identify fairness reviews from a large number of reviews. However, they do not provide a fine-grained view of different types of fairness concerns raised by various users of AI-based apps. Hence, we aimed to cluster and summarize fairness reviews, as discussed next.
@@ 5.1.1. Dataset
As discussed in Section 4.2.1, Logistic Regression was selected as the best-performing model because it had the highest precision and AUC. Once we applied it on 9,475,506 reviews, we got 91,974 fairness reviews (see Fig. 4). We used these 91,974 reviews as a dataset for the clustering and summarization approach.
@@ 5.1.2. K-means Clustering
In addition to the summarization metric, we used the silhouette score to determine the quality of clusters (Caliński and Harabasz, 1974). The silhouette score is a measure of how similar a data point is to the other data points in its cluster compared to how similar it is to the data points in other clusters. A low silhouette score of a review implies that the review could be equally well assigned to another cluster. The silhouette score thus indicates whether a given cluster is well-formed or not. We executed K-means with different $k$ ranging from 2 to 10. We did not go beyond 10 as our observations in the labeling process (see Section 3.1) provided clues that there should not be more than 10 dominant fairness concerns in user reviews. However, as discussed in (Nema et al., 2022), a larger $k$ could have helped us identify the long-tail of fairness concerns. For each $k$, K-means was run 100 times with different centroid seeds. We also set the maximum number of iterations for a single run to 100. The final result of K-means is the best output out of 100 runs in terms of inertia, which is the sum of squared distances of samples to their closest cluster center. Our experiments showed that the highest value for the summarization metric is achieved when k=10 (Fig. 4). Fig. 5 presents the distribution of the silhouette scores of reviews in each of the 10 clusters (k=10). The dotted red line shows the average silhouette score for k=10. As shown in Fig. 5, at least 40% of the reviews in clusters 2 to 9 have a silhouette score more than the average (the dotted red line). Clusters 2 to 9 are considered compact clusters. Clusters 0 and 1 were excluded from further analysis because they did not demonstrate a high degree of cohesion, as more than 60% of the reviews in these clusters had a silhouette score below the average.
@@
To understand the main (dominant) fairness concern discussed in each of the eight compact clusters, we descendingly sorted the reviews in each compact cluster based on their silhouette scores. The first author manually analyzed the 30 reviews with the highest silhouette scores in each cluster ($8*30=240$ reviews in total) and suggested a topic for each cluster. Then, three other authors rechecked all these 240 reviews and the proposed topics. The first author and the rest of the authors held several meetings to discuss the topics, refine them, and solve any disagreements. Although we had eight compact clusters, as shown in Fig. 4, the qualitative analysis process led to the identification of six distinct, dominant fairness concerns that are discussed next in Section 5.2. This was achieved by merging the topics emerging from some clusters that exhibited overlapping concerns. This approach was taken to provide a more concise and distinct set of fairness concerns. For example, we identified two clusters with approximately overlapping topics. One highlighted user concerns about ‘disparities in feature quality and service access across platforms and devices’, and another included complaints about ‘being deprived of some features and services in platforms and devices’. Given the closeness of these concerns, we merged them into a single fairness concern called ‘receiving different quality of features and services in different platforms and devices’. This new fairness concern encapsulates the essence of both initial fairness concerns. Merging fairness concerns was done thoughtfully to ensure that the distinct nature of each concern was preserved while avoiding redundancy.
@@ 5.2. Results
6. Root Causes of Fairness Concerns (RQ3)
@@ 6.1. Approach
As shown in Fig. 4, our model identified 91,974 fairness reviews (Section 4.2.1), which were clustered into six types of fairness concerns (Section 5.2). These 91,974 app reviews received 4,911 responses from app owners. To understand the root causes reported by app owners for the identified fairness concerns, we randomly selected 2,248 out of the 4,911 app owners’ responses to fairness reviews with a confidence level of 99% and margin error of 2% (Israel et al., 1992; Kadam and Bhalerao, 2010). In the first step, the first author conducted the open coding procedure (Glaser and Strauss, 2017) on these 2,248 app owners’ responses. Next, the third author reviewed all the codes and categories produced by the first author. Last, the first and third authors held several meetings to discuss and refine the codes and categories. Similar to Section 3.1, the negotiated agreement approach (Campbell et al., 2013) was employed to resolve any disagreements and conflicts between the authors. This process led to the identification of six root causes of fairness concerns.
@@ 6.2. Results
1. Copyright Issues. App owners admit that some of the fairness concerns are due to the copyright issues imposed on the apps, platforms, or devices. This can mainly be a reason behind Fairness Concern 1 (receiving different quality of features and services in different platforms and devices) identified in RQ2 (Section 5.2). For example in the following conversation, the copyright issue is mentioned by an app owner as the reason why Android and iOS users do not receive the same emojis.
@@
3. Faults. Some app owners admit that faults in AI-based apps could contribute to the identified fairness concerns (particularly Fairness Concerns 4 and 5). For example, in the following response, the owner accepted that an error may have led to incorrectly banning the user.
@@
Table 6. Frequency of fairness concerns (clusters) in each category. FC: Fairness Concern
Categories FC1 FC2 FC3 FC4 FC5 FC6
Photography 492 63 84 571 95 442
Productivity 417 286 28 334 93 304
Communication 11,942 716 576 2,097 550 362
Shopping 27 146 60 308 112 302
Tools 1,247 1,222 107 500 862 121
Social 2,969 1,794 5,706 3,922 3,904 382
Video Players & Editors 182 313 3,656 572 1,880 1,555
Travel & Local 35 82 60 229 197 240
Entertainment 250 89 20 239 115 158
Finance 8 1 2 29 8 139
Business 5 11 1 45 9 49
Music & Audio 210 86 33 364 788 1,889
Health & Fitness 22 12 9 165 88 283
Food & Drink 217 6 1 150 51 123
Lifestyle 21 10 1 37 2 47
Role Playing 4 6 17 20 15 607
Arcade 0 2 2 32 2 175
Education 18 1 0 26 1 101
Strategy 3 3 3 11 2 107
Art & Design 15 1 1 61 0 28
Auto & Vehicles 0 0 1 1 0 1
Puzzle 1 0 0 0 0 2
Personalization 5 0 0 2 0 1
Total 18,090 4,850 10,368 9,715 8,774 7,418
@@ 7. Discussion
Benefits of findings for mobile app development tasks. In a systematic review, Dąbrowski et al. (Dąbrowski et al., 2022) found that app review analysis can support various software development activities (e.g., requirement engineering and design). We argue that the findings obtained in RQ2 (fairness concerns from the users’ perceptive) and RQ3 (root causes of fairness concerns from the app owners’ viewpoint) could benefit mobile app development in different ways. (1) The identified fairness concerns and their potential root causes can help app owners and developers understand and elicit users’ fairness-related requirements and expectations. For example, the identified fairness concerns can guide app developers on questions that should be asked of users to gather fairness-related requirements and expectations. (2) The knowledge of potential fairness concerns can help app developers be aware of and avoid common fairness concerns during app development. (3) The identified fairness concerns can act as a checklist for app developers to enable them to specify test scenarios that evaluate specific fairness concerns. (4) Our findings can inform researchers and tool builders to develop tools to embed fairness in AI-based apps, and detect and fix fairness concerns. (5) The six identified fairness concerns could serve as the basis for constructing multi-label datasets of fairness concerns and developing multi-classification techniques that automatically detect these fairness concerns in text-based software artifacts.
@@ 8. Threats to Validity
Internal Validity. A key concern in our internal validity is the inherent ambiguity attached to the concept of ‘fairness’, which can be nuanced, with different interpretations. App users might express fairness concerns using various terms and phrases, not all of which are easy to capture by a keyword-based approach like ours. To mitigate the potential impact of this threat, we used an iterative and multi-faceted approach to identify fairness-related keywords. We started with a literature-driven set of keywords and expanded this using KeyBERT. The keyword set was reviewed and refined manually to ensure relevance and accuracy. Moreover, our manual labeling process in RQ1 and qualitative analysis in RQ2 and RQ3 included multiple coders to minimize individual bias and capture a broader interpretation of fairness.
Conclusion Validity. The conclusions regarding the fairness concerns of users and the root causes of the fairness concerns were based on the results of our clustering approach. As discussed in Section 5, we excluded Clusters 0 and 1 as they were not compact clusters. These two clusters together contained around 18,090 reviews. Such exclusion could have led to missing some important types of fairness concerns. Furthermore, we only experimented with $k$=2 to 10 to detect the optimum number of clusters. We acknowledge that investigating a larger $k$ may lead to identifying more or different fairness concerns, particularly the long-tail of fairness concerns. We also concede that different instrumentation of our clustering approach might have influenced our conclusions in Section 5.2 and Section 6.2. To mitigate the potential impact of this threat, we strived for transparency and thoroughness in our reporting. We explained our choices in the clustering process, including why a particular technique was selected and our parameter-tuning process. We also acknowledge the influence of these choices on our findings. While our choices were reasonable and based on best practices, further experimentation with more techniques is required in our future work. Finally, Table 1 illustrates that 7 categories had only one representative app in our analysis. This may have impacted our conclusions on fairness concerns in app categories (see Table 6). Future research could build upon our work by expanding our dataset to include a larger and equal number of apps in each category, enhancing the findings’ reliability.
External Validity. Our evaluation is based on 108 apps from 23 categories and an overall analysis of $\approx$9.5 million reviews. The results we received across these apps with a thorough manual analysis of $>$2,600 reviews for building our model, analysis of $>$92,000 reviews using the model and clustering, and qualitative analysis of 2,248 lend a degree of confidence to the generalizability of our study. Further experimentation is nonetheless required to mitigate external validity threats.